# Dissipative qutrit-mediated stable charging

### v1

@@ Abstract
In this work, we propose a stable charging scheme mediated by a dissipative three-level system, which allows a unidirectional energy flow from the external power source to a finite-dimensional quantum battery. By virtue of the dissipation, the battery avoids the spontaneous discharging induced by the time-reversal symmetry in any unitary charging scheme. Irrespective to the initial state, the battery can be eventually stabilized at the maximal-ergotropy state as long as the charger-battery interaction is present. We use a Dyson series of Lindbladian superoperator to obtain an effective master equation to describe the battery dynamics and extract the optimization condition for charging efficiency. Also it proves that the Fermi’s golden rule can adapt to the non-Hermitian Hamiltonian. We justify the outstanding performance of our charging scheme when applied to the finite-size system with a uniform energy ladder operator, the large spin battery, and the truncated harmonic oscillator battery.
@@ I Introduction
As a microscopic unit for energy storage, quantum battery Campaioli et al. (2024a) has recently received remarkable attention as both a promising energy supply device for quantum technology Auffèves (2022) and a tool to develop the wide application of quantum thermodynamics Horodecki and Oppenheim (2013); Hovhannisyan et al. (2013); Vinjanampathy and Anders (2016); Campaioli et al. (2018). Since formally introduced by Alicki and Fannes Alicki and Fannes (2013), a lot of efforts have been devoted to studying the role of quantum resources such as quantum correlation and quantum coherence in the charging performance of various quantum batteries Alicki and Fannes (2013); Campaioli et al. (2017); Kamin et al. (2020); Andolina et al. (2019); Seah et al. (2021); Arjmandi et al. (2022a); Shi et al. (2022); Yang et al. (2024). It is found that the collective charging schemes could achieve a super-extensive scaling in power, which supports that the quantum correlation provides a dramatic boost to the charging efficiency of the parallel charging schemes Ferraro et al. (2018); Le et al. (2018); Rossini et al. (2020). Quantum speed up in charging can also be found in a single high-dimensional quantum battery via repeated collisions between chargers and battery Seah et al. (2021); Salvia et al. (2023); Beleño et al. (2024). In addition, quantum measurements on chargers can give rise to a more efficient charging performance than the measurement-free schemes and do not invoke the initial quantum coherence in both battery and chargers Yan and Jing (2023).
@@
Numerous prototypes of quantum batteries have been developed in diverse models. The one-dimensional spin chain battery has been proposed in Ref. Le et al. (2018). Then many investigations further extended the spin chain to the spin-ensemble systems Qi and Jing (2021); Huangfu and Jing (2021); Zhao et al. (2021); Arjmandi et al. (2022b); Barra et al. (2022); Gao et al. (2022); Ghosh and Sen (De); Yao and Shao (2022). For example, by mediating between the charger and battery both consisting of a spin ensemble through a magnon system, we can realize a long-range charging Qi and Jing (2021). Dick model battery is another type of extensively studied model Ferraro et al. (2018); Andolina et al. (2019); Zhang et al. (2019); Crescente et al. (2020a, b), which is based on the interaction between an ensemble of noninteracting two-level systems (TLSs) and a common cavity mode. The connection and distinction between the spin-charger and cavity-charger schemes were discussed in Ref. Huangfu and Jing (2021). In addition to the spin systems, the bosonic quantum battery based on an ensemble of harmonic oscillators (HOs) was proposed in Ref. Friis and Huber (2018) focusing on the charging precision, i.e., the variance of high-level population distribution.
@@
Besides the charging efficiency, stabilizing a quantum battery at the charged (high-ergotropy) state is also crucial for the reliable operation over quantum batteries. Many conventional schemes have to immediately decouple the charger from the battery of a finite size when the latter reaches the fully charged state. If they remain coupled with each other, energy can flow back from the battery to the charger, degrading the charging-efficiency and the energy retention Ferraro et al. (2018); Zhang et al. (2019); Andolina et al. (2018, 2019); Guo et al. (2024). Extra operations have been designed to address this challenge. The Zeno-protection scheme applies sequential projective measurements on the high energy state of a two-level quantum battery Gherardini et al. (2020). The time-dependent Hamiltonian schemes use the stimulated Raman adiabatic passage to maintain a stable charged state Santos et al. (2019); Rojo-Francàs et al. (2024). Nonlocal and strong chaotic correlation was demonstrated in a Sachdev-Ye-Kitaev quantum battery Rosa et al. (2020) to improve the charging stability, to name a few.
@@
An alternative approach for stable charging is assisted by a dissipative charger connecting the battery and the external power source Farina et al. (2019). In that model that both charger and battery are identical HOs or TLSs, the energy and ergotropy of the battery can be stabilized at a finite value determined by the charger-battery coupling strength and the intensity of the external power source. However, the battery could not be stabilized at its highest energy level in such a setting. Introducing a common dissipative environment for the composite system of battery and charger and carefully designing their inner coupling strength and the dissipative rates can increase the steady energy in the battery Ahmadi et al. (2024).
@@
In this work, we introduce an efficient and stable charging model that the charger is modeled as a dissipative $\Delta$-type qutrit mediating between the external power source and the battery. The charger is coupled to the battery through the energy exchange interaction. By deriving an effective master equation for the battery using a perturbation technique for Lindbliadian superoperator, we find that a unidirectional energy flow from the external source to the battery is constructed due to the time-reversal symmetry breaking by the dissipative charger. We first consider the battery with a uniform energy ladder operator that is popular in many worksSeah et al. (2021); Chen et al. (2022); Mitchison et al. (2021). Our scheme is characterized with universality in enhancing and holding the battery energy for any initial state Sathe and Caravelli (2024) and does not require the charger to be decoupled from the battery after the charging is completed. We apply our scheme to more realistic scenarios by modeling quantum battery as a large spin and a truncated quantum harmonic oscillator. It is interesting to find that choosing an appropriate charger-battery coupling strength or adjusting it during the charging process can significantly boost the charging efficiency.
@@
The rest of the paper is organized as follows. In Sec. II, we introduce our dissipative-mediated charging model. In Sec. III, we derive an effective master equation for the battery in a weak-driving regime (see appendix A for details) and obtain the optimization condition for charging efficiency. In a broader perspective, we actually develop an efficient technique to extend the Fermi’s Golden rule to adapt to the non-Hermitian Hamiltonian. In Sec. IV, we discuss the large spin quantum battery. It is found that charging efficiency can be significantly boosted by optimizing the effective transition rate for special subspaces. In Sec. V, we discuss the truncated harmonic oscillator quantum battery. It is shown that by adjusting the charger-battery coupling strength according to the instantaneous state of the battery, the charging efficiency can be greatly enhanced. The conclusion is provided in Sec. VI.
@@ II Model
Our indirect charging model consists of a quantum battery of a finite size and a dissipative charger that is driven by an external power source. As shown in Fig. 1, the charger is a dissipative $\Delta$-type qutrit. An external field directly drives the transition between levels $|g\rangle$ and $|h\rangle$. The quantum battery in Secs. II and III has $N+1$ uniform energy levels with a bare Hamiltonian ($\hbar=1$) $H_{B}=E\sum_{n=0}^{N}n|n\rangle\langle n|$. The charging process is switched on when the quantum battery couples to the qutrit via the transition $|e\rangle\leftrightarrow|h\rangle$. The full system Hamiltonian reads,
@@
$U=\exp\left\{i\left[H_{B}+\omega_{L}|h\rangle\langle h|+(\omega_{L}-E)|e\rangle\langle e|\right]t\right\},$ (2)
@@
$H^{\prime}=$ $UHU^{\dagger}-iU\frac{\partial}{\partial t}U^{\dagger}=\Delta|h\rangle\langle h|+\delta|e\rangle\langle e|$ (3)
@@
with detunings $\Delta=\omega_{hg}-\omega_{L}$ and $\delta=\omega_{eg}+E-\omega_{L}$. When the charger is subject to a local dissipative environment, the evolution of the charger-battery system is governed by the master equation Carmichael (1999):
@@ III Effective master equation
Figure 2: (a) Main mechanism of our charging scheme. The external power source $\Omega$, the charger-battery interaction $gA_{n}$, and the spontaneous decay $\gamma_{eg}$ induces the transitions $|gn\rangle\rightarrow|hn\rangle$, $|hn\rangle\rightarrow|e(n+1)\rangle$, and $|e(n+1)\rangle\rightarrow|g(n+1)\rangle$, respectively. They give rise to an effective transition $|gn\rangle\rightarrow|g(n+1)\rangle$ with a rate $\gamma_{\rm eff}$. (b) Leading-order transitions induced by the atomic spontaneous emissions from $|h\rangle$ to $|e\rangle$ and from $|e\rangle$ to $|g\rangle$. The decay channel indicated by $\gamma_{he}$ is the main negative factor for charging.
In our charging scheme, the mediator qutrit is initially prepared in its ground state $|g\rangle$. In the weak-driving regime $\Omega\ll\Delta,\ \gamma_{hg}$, one can suppose that the qutrit subject to a dissipative environment remains mostly in its ground state $\rho(t)\approx P_{g}\rho(t)P_{g}$, where $P_{g}\equiv|g\rangle\langle g|$. According to the master equation (4), the main mechanism underlying the indirect charging can be illustrated in the subspaces spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ for $n<N$. As shown in Fig. 2(a), the system is driven from $|gn\rangle$ to $|hn\rangle$ through the external power source $\Omega$ and from $|hn\rangle$ to $|e(n+1)\rangle$ via the charger-battery interaction with a coupling strength $gA_{n}$. Subsequently, the spontaneous decay $\gamma_{eg}$ induces a transition from $|e(n+1)\rangle$ to $|g(n+1)\rangle$. These processes constitute an effective unidirectional transition from $|gn\rangle$ to $|g(n+1)\rangle$, which can be quantitatively described by the effective Lindblad operator in Eq. (48):
@@
with $\tilde{\Delta}\equiv\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}\equiv\delta-i\gamma_{eg}/2$ the complex detunings. Due to the large-detuning condition, the decay from $|hn\rangle$ to $|gn\rangle$ contributes to a pure dephasing of the state $|gn\rangle$ in the leading order. The relevant effective Lindblad operator given by Eq. (47) reads
$L_{hg,n}^{\rm eff}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|.$ (6)
In addition, the energy shift of $|gn\rangle$ induced by the weak driving and the charger-battery interaction can be described by the effective Hamiltonian
@@
By tracing out the qutrit from Eq. (4), the effective equation of motion of the battery is given by
@@
$L^{B}_{eg,{\rm eff}}$ $=\sum_{n=0}^{N}\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|n+1\rangle\langle n|=\frac{\sqrt{\gamma_{eg}}\Omega}{A^{\dagger}Ag^{2}-\tilde{\Delta}\tilde{\delta}}A^{\dagger}.$ (10b)
are the effective Hamiltonian and Lindblad operators in the battery Hilbert space, respectively. When expanding the coefficients of the effective Lindblad operators in terms of series
$\langle gn|L_{hg,n}^{\rm eff}|gn\rangle$ $=\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{\Omega}{\tilde{\Delta}}$ (11)
$=-\sqrt{\gamma_{hg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k)},$
$\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle$ $=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\frac{gA_{n}}{\tilde{\delta}}\frac{\Omega}{\tilde{\Delta}}$
$=-\sqrt{\gamma_{eg}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)},$
it is interesting to find that $\Omega_{\rm eff}^{(l)}$’s are precisely the effective coupling strengths connecting $|i\rangle=|gn\rangle$ and $|f\rangle=|hn\rangle$ or $|e(n+1)\rangle$ up to the first order in $\Omega$ and $l$ order in $gA_{n}$ with respect to the non-Hermitian Hamiltonian in Eq. (31):
$\tilde{H}_{n}$ $=\tilde{H}_{0,n}+\tilde{V}_{n},$ (12)
@@
$\tilde{V}_{n}$ $=gA_{n}(|hn\rangle\langle e(n+1)|+|e(n+1)\rangle\langle hn|).$
Particularly, they can be obtained by the constrained high-order Fermi’s golden rules Kockum et al. (2017) with the form
@@
where $\tilde{V}_{n,m_{i}m_{j}}=\langle m_{i}|\tilde{V}_{n}|m_{j}\rangle$ with $m_{i}=hn,e(n+1)$ the intermediate states. $\tilde{E}_{m_{i}}=\langle m_{i}|\tilde{H}_{0,n}|m_{i}\rangle$ and $\tilde{E}_{i}=\langle gn|\tilde{H}_{0,n}|gn\rangle=0$ are the diagonal elements of $\tilde{H}_{0,n}$. For an even $l$, $\Omega_{\rm eff}^{(l)}$ represents the effective coupling strength for connecting $|gn\rangle$ and $|hn\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|hn\rangle$. For an odd $l$, it represents the effective coupling strength for connecting $|gn\rangle$ and $|e(n+1)\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|e(n+1)\rangle$. Similarly, the coefficient of the effective Hamiltonian (9) is the real part of the summation of all round trips about $|gn\rangle$.
@@
where $\phi=\arg(\tilde{\Delta}\tilde{\delta})$. With fixed detunings $\tilde{\Delta}$ and $\tilde{\delta}$, one can find that the effective transition rate in the $n$th subspace can be optimized by setting
$g^{2}_{\rm opt}A^{2}_{n}=|\tilde{\Delta}\tilde{\delta}|.$ (15)
The decay channel $|h\rangle\rightarrow|e\rangle$ plays a negative role in the charging efficiency through a higher-order process than $|e\rangle\rightarrow|g\rangle$ and $|h\rangle\rightarrow|g\rangle$ (see Appendix A). As the leading-order contribution depicted in Fig. 2(b), the system in the subspace is driven by the external power from $|gn\rangle$ to $|hn\rangle$ then decay from $|hn\rangle$ to $|en\rangle$ and then decay from $|en\rangle$ to $|gn\rangle$ will give rise to an effective dephasing on the state $|gn\rangle$ with a rate given by Eq. (58). In a biased fluxonium system at about $0.05\leq\Phi_{\rm ex}/\Phi_{0}\leq 0.15$, where $\Phi_{\rm ex}$ and $\Phi_{0}$ are the applied flux and flux quantum, respectively, $\gamma_{he}$ is the same order in magnitude as $\gamma_{eg}$ and is one order of magnitude smaller than $\gamma_{hg}$ Joo et al. (2010); Manucharyan et al. (2009), i.e. $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$. With such a parametric setting, one can confirm that the dephasing rate caused by the process in Fig. 2(b) is negligible in comparison to that described by Eq. (6) for $|gn\rangle\rightarrow|hn\rangle\rightarrow|gn\rangle$ in Fig. 2(a). Similarly, the process that from $|gn\rangle$ to $|hn\rangle$ through the external driving, from $|hn\rangle$ to $|en\rangle$ through the spontaneous decay $\gamma_{he}$, from $|en\rangle$ to $|h(n-1)\rangle$ through the charger-battery interaction, and finally from $|h(n-1)\rangle$ to $|g(n-1)\rangle$ through the spontaneous decay $\gamma_{hg}$ gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$ with a rate given by Eq. (61). It is also negligible comparing to the effective transition rate in Eq. (14) under the condition $|\tilde{\Delta}|\gg|\tilde{\delta}|$ and $\gamma_{eg}\approx\gamma_{he}$. Consequently, the presence of a decay channel $|h\rangle\rightarrow|e\rangle$ does not significantly alter our unidirectional energy flow.
Figure 3: Comparison between the dynamics obtained by the full master equation (4) and the effective master equation (8) of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with various $\Omega$. The coupling strength $g$ satisfies the optimized condition (15). The battery size is set as $N=50$. $\gamma_{he}=0$, $\gamma_{hg}=10\gamma_{eg}=0.1E_{B}$, and $\Delta=10\delta=0.1E_{B}$.
Figure 4: Charging performance in the presence of the decay from $|h\rangle$ to $|e\rangle$ with $\gamma_{he}=\gamma_{eg}$ under various inverse temperature $\beta$ ($k_{B}\equiv 1$). (a) Energy stored in the battery $\Delta E(t)$. (b) Ergotropy $\mathcal{E}(t)$. The driving intensity is fixed as $\Omega=0.05\gamma_{hg}$. Other parameters are the same as in Fig. 3.
The justification of our charging scheme begins with a simple case with $A_{n}=1$, resulting in uniform effective transition rates across all the subspaces. Substituting the optimized condition (15) to Eq. (14), it is found that the optimized effective transition rate reads

### v2

@@ Abstract
In this work, we propose a stable charging scheme mediated by a three-level system (qutrit), which renders a unidirectional energy flow from an external power source to an $(N+1)$-dimensional quantum battery. By virtue of the qutrit dissipation, the battery avoids the spontaneous discharging induced by the time-reversal symmetry of any unitary-charging scheme. Irrespective of the initial state, the battery can be eventually stabilized at the maximal-ergotropy state as long as the charger-battery interaction is present. We use a Dyson series of Lindbladian superoperator to obtain an effective master equation for the battery, which is found to be equivalent to the high-order Fermi’s golden rule adapted to the non-Hermitian Hamiltonian and spontaneous decay. We extract the optimization condition for charging efficiency and justify it in the finite-size battery with uniform energy splitting, the large spin battery, and the truncated harmonic-oscillator battery.
@@ I Introduction
Quantum battery is a microscopic unit for storing energy Campaioli et al. (2024a). Recently, it has received remarkable attention as a promising energy supply device for quantum technology Auffèves (2022) and a tool to develop a wide range of applications of quantum thermodynamics Horodecki and Oppenheim (2013); Hovhannisyan et al. (2013); Vinjanampathy and Anders (2016); Campaioli et al. (2018). Since formally introduced by Alicki and Fannes Alicki and Fannes (2013), much effort has been devoted to studying the role of quantum resources, e.g., correlation, coherence, and entanglement, in the charging performance of various quantum batteries Alicki and Fannes (2013); Campaioli et al. (2017); Kamin et al. (2020); Andolina et al. (2019); Seah et al. (2021); Arjmandi et al. (2022a); Shi et al. (2022); Yang et al. (2024). The collective charging schemes could achieve a superextensive scaling in power Ferraro et al. (2018); Le et al. (2018); Rossini et al. (2020). It is demonstrated that the quantum correlation provides a dramatic boost to the charging efficiency of the parallel schemes. Repeated collisions between chargers and battery speeds up the charging rate of a finite-dimensional quantum battery Seah et al. (2021); Salvia et al. (2023); Beleño et al. (2024). Quantum measurements on chargers could give rise to a more efficient charging performance than the measurement-free schemes and do not take advantage of the initial coherence in both battery and chargers Yan and Jing (2023). In addition, an implementation-independent approach indicates that the charging power is not an entanglement monotone Gyhm and Fischer (2024).
@@
Numerous prototypes of quantum batteries have been developed in diverse models. A one-dimensional spin chain battery has been proposed in Ref. Le et al. (2018). Afterwards many investigations extended the spin chain Grazi et al. (2025) to the spin-ensemble systems Qi and Jing (2021); Huangfu and Jing (2021); Zhao et al. (2021); Arjmandi et al. (2022b); Barra et al. (2022); Gao et al. (2022); Ghosh and Sen (De); Yao and Shao (2022). For example, by mediating between the charger and battery both consisting of a spin ensemble through a magnon system, one can realize a long-range and full charging Qi and Jing (2021). Dicke-model battery is also under extensive studies Ferraro et al. (2018); Andolina et al. (2019); Zhang et al. (2019); Crescente et al. (2020a, b), which is based on the interaction between an ensemble of noninteracting two-level systems (TLSs) and a common cavity mode. The similarity and distinction between the spin-charger and cavity-charger schemes were discussed in Ref. Huangfu and Jing (2021). In addition to the spin systems, the bosonic quantum battery based on an ensemble of harmonic oscillators (HOs) was proposed in Ref. Friis and Huber (2018), focusing on the charging precision, i.e., the variance of high-level population distribution.
@@
Besides the charging efficiency, stabilizing the battery system at the highly charged (high-ergotropy) state is also crucial for reliable quantum batteries. Many conventional schemes have to immediately decouple the charger from the battery when the latter is fully charged. Otherwise, the stored energy can flow back from the battery to the charger, degrading the charging performance and the energy retention Ferraro et al. (2018); Zhang et al. (2019); Andolina et al. (2018, 2019); Guo et al. (2024). Extra operations have been designed to address discharging. Zeno-protection scheme applies sequential projective measurements on the upper state of a two-level quantum battery Gherardini et al. (2020). Selective weak measurements were implemented to protect the battery from discharging by decoherence without extra net recharging Malavazi et al. (2025). The stimulated Raman adiabatic passage was used to maintain a stable charged state Santos et al. (2019); Rojo-Francàs et al. (2024). Nonlocal and strong chaotic correlation was demonstrated in a Sachdev-Ye-Kitaev quantum battery Rosa et al. (2020) to improve the charging stability.
@@
An alternative approach for stable charging is assisted by a dissipative charger connecting the battery and the external power source Farina et al. (2019), where both charger and battery are identical HOs or TLSs. The energy and ergotropy of the battery can be stabilized with a certain amount determined by the charger-battery coupling strength and the intensity of the external power source. However, the battery could not be maintained at its highest energy level in such a setting. Introducing a common dissipative environment for battery and charger might induce a unidirectional energy flow and increase the steady energy in the HO battery Ahmadi et al. (2024). While that scheme was much sensitive to the subtle balance between the charger-battery coupling strength and their dissipative rates.
@@
In this work, we introduce an efficient and stable charging model that the charger is modeled as a dissipative $\Delta$-type qutrit mediating between the external power source and the battery. The charger is coupled to the battery through the exchange interaction. By deriving an effective master equation about the battery using a perturbation technique for Lindbladian superoperator, we find that a unidirectional energy flow from the external source to the battery is constructed due to the time-reversal symmetry broken by the dissipative charger. Our scheme is universal in enhancing and holding the battery energy for any initial state Sathe and Caravelli (2025) and needs not decouple the charger from the battery after the charging is completed. We apply it to three battery prototypes with various coupling strengths. It is interesting to find that choosing an appropriate charger-battery coupling strength or quenching it during the charging process can significantly boost the charging efficiency.
@@
The rest of the paper is organized as follows. In Sec. II, we introduce our dissipative-mediated charging model. In Sec. III, we derive an effective master equation for the battery in a weak-driving regime (see appendix A for details) and obtain the optimization condition for charging the uniform-splitting battery. In a broader perspective, we develop an efficient technique to extend the Fermi’s golden rule to the situation with non-Hermitian Hamiltonian (battery) and spontaneous decay (charger). In Sec. IV, we discuss the large-spin battery. It is found that charging efficiency can be significantly boosted by optimizing the effective transition rate for special subspaces. In Sec. V, we discuss the truncated harmonic-oscillator battery. It is shown that by adjusting the charger-battery coupling strength due to the instantaneous state of the battery, the charging efficiency can be greatly enhanced. The conclusion is provided in Sec. VI.
@@ II Model
Our indirect charging model consists of a quantum battery and a charger that is driven by an external power source. As depicted by Fig. 1, the quantum battery in Secs. II and III has $N+1$ uniform energy levels with a bare Hamiltonian ($\hbar=1$) $H_{B}=E_{B}\sum_{n=0}^{N}n|n\rangle\langle n|$. The charger is a dissipative $\Delta$-type qutrit. An external field directly drives the transition between levels $|g\rangle$ and $|h\rangle$. The charging process is switched on when the quantum battery couples to the qutrit via the transition $|e\rangle\leftrightarrow|h\rangle$. The full system Hamiltonian reads,
@@
$U=\exp\left\{i\left[H_{B}+\omega_{L}|h\rangle\langle h|+(\omega_{L}-E_{B})|e\rangle\langle e|\right]t\right\},$ (2)
@@
$H^{\prime}=$ $UHU^{\dagger}-iU\frac{\partial U^{\dagger}}{\partial t}=\Delta|h\rangle\langle h|+\delta|e\rangle\langle e|$ (3)
@@
with detunings $\Delta=\omega_{hg}-\omega_{L}$ and $\delta=\omega_{eg}+E_{B}-\omega_{L}$. When the charger is subject to a local dissipative environment, the evolution of the charger-battery system is governed by the master equation Carmichael (1999):
@@ III Effective master equation
Figure 2: (a) Main mechanism of our charging scheme. The external power source $\Omega$, the charger-battery interaction $gA_{n}$, and the spontaneous decay $\gamma_{eg}$ induces the transitions $|gn\rangle\leftrightarrow|hn\rangle$, $|hn\rangle\leftrightarrow|e(n+1)\rangle$, and $|e(n+1)\rangle\rightarrow|g(n+1)\rangle$, respectively. They give rise to an effective transition $|gn\rangle\rightarrow|g(n+1)\rangle$ with a rate $\gamma_{n,{\rm eff}}$. (b) Leading-order discharging transition induced by the atomic spontaneous emissions from $|hn\rangle$ to $|en\rangle$ and from $|h(n-1)\rangle$ to $|g(n-1)\rangle$. The decay channel indicated by $\gamma_{he}$ is the main negative factor for charging.
In our charging scheme, the mediator qutrit is initially prepared in its ground state $|g\rangle$. In the weak-driving regime, i.e., $\Omega<\gamma_{eg}$ and $\Omega\ll\Delta$, the qutrit subject to a dissipative environment can be supposed to remain mostly in its ground state $\rho(t)\approx P_{g}\rho(t)P_{g}$, where $P_{g}\equiv|g\rangle\langle g|$. According to the full master equation (4), the main mechanism underlying the indirect charging can be illustrated in the subspaces spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ for $n<N$. Figure 2(a) describes the leading-order contribution from the driving Hamiltonian $\Omega(|h\rangle\langle g|+|g\rangle\langle h|)$, the coupling Hamiltonian $g(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|)$, and the spontaneous decay $\mathcal{L}[\sqrt{\gamma_{eg}}|g\rangle\langle e|]$. In another word, the three-step process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\rightarrow|g(n+1)\rangle$ and its correction of higher orders by repeating $|hn\rangle\leftrightarrow|e(n+1)\rangle$ constitute the charging mechanism: $|gn\rangle\rightarrow|g(n+1)\rangle$. Meanwhile, the leading-order contribution from the driving $\Omega(|h\rangle\langle g|+|g\rangle\langle h|)$, the decay $\mathcal{L}[\sqrt{\gamma_{he}}|e\rangle\langle h|]$, the charger-battery coupling $g(A^{\dagger}|e\rangle\langle h|+A|h\rangle\langle e|)$, and the decay $\mathcal{L}[\sqrt{\gamma_{hg}}|g\rangle\langle h|]$ is described by Fig. 2(b). Specifically, the four-step process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ and its correction of higher orders by repeating $|en\rangle\leftrightarrow|h(n-1)\rangle$ constitute the discharging mechanism: $|gn\rangle\rightarrow|g(n-1)\rangle$. In regard of the perturbation theory, the discharging process is one order higher than the charging process and then charging dominates discharging. This nonreciprocity results in a unidirectional energy flow from the external source to the battery.
Considering Fig. 2(a) for the $n$th subspace, the charging processes can be quantitatively described by the effective Lindblad operator in Eq. (48):
@@
with $\tilde{\Delta}\equiv\Delta-i\gamma_{hg}/2-i\gamma_{he}/2$ and $\tilde{\delta}\equiv\delta-i\gamma_{eg}/2$ the complex detunings. Due to the large-detuning condition, the spontaneous decay from $|hn\rangle$ to $|gn\rangle$ contributes to a pure dephasing of the state $|gn\rangle$ in the leading order. The relevant effective Lindblad operator reads
$L_{hg,n}^{\rm eff}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|$ (6)
as given by Eq. (47). In addition, the energy shift of $|gn\rangle$ induced by the weak driving and the charger-battery interaction can be described by the effective Hamiltonian
@@
By tracing out the qutrit from the full master equation (4), the effective equation of motion for the battery under charging is given by
@@
$L^{B}_{eg,{\rm eff}}$ $=\sum_{n=0}^{N-1}\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|n+1\rangle\langle n|=\frac{\sqrt{\gamma_{eg}}\Omega}{A^{\dagger}Ag^{2}-\tilde{\Delta}\tilde{\delta}}A^{\dagger}.$ (10b)
are the effective Hamiltonian and Lindblad operators in the whole Hilbert space of battery, respectively.
Through expanding the coefficients of the effective Lindblad operators in the $n$th subspace:
$\langle gn|L_{hg,n}^{\rm eff}|gn\rangle$ $=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}$
$=\frac{\sqrt{\gamma_{hg}}}{\tilde{\Delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)},$ (11a)
$\langle g(n+1)|L_{eg,n}^{\rm eff}|gn\rangle$ $=-\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{\tilde{\Delta}\tilde{\delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}$
$=\frac{\sqrt{\gamma_{eg}}}{\tilde{\delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+2)},$ (11b)
it is interesting to find that $\Omega_{\rm eff}^{(l)}$’s are precisely the effective coupling strengths linking $|i\rangle=|gn\rangle$ and $|f\rangle=|hn\rangle$ or $|e(n+1)\rangle$ up to the first order in $\Omega$ and (l1)(l-1) order in $gA_{n}$ with respect to the non-Hermitian Hamiltonian $H_{\rm NH}$ in Eq. (27). In the $n$th subspace as shown in Fig. 2(a), we have
$(H_{\rm NH})_{n}$ $=\tilde{H}_{0,n}+\tilde{V}_{n},$ (12)
@@
$\tilde{V}_{n}$ $=gA_{n}|hn\rangle\langle e(n+1)|+\Omega|hn\rangle\langle gn|+{\rm H.c.}$
For a standard perturbative derivation, $(H_{\rm NH})_{n}$ in appendix A is partitioned into the excited-subspace Hamiltonian $\tilde{H}_{n}$ and the weak driving Hamiltonian $V_{n}$, given by Eqs. (31) and (32), respectively. In Eq. (12), it is divided into the diagonal component $H_{0,n}$ and the off-diagonal one $\tilde{V}_{n}$. Particularly, $\Omega_{\rm eff}^{(l)}$ can be obtained by the constrained Fermi’s golden rules Kockum et al. (2017) in the form
@@
where $\tilde{V}_{n,m_{i}m_{j}}=\langle m_{i}|\tilde{V}_{n}|m_{j}\rangle$ with $m_{i}=hn,e(n+1)$ the excited states, $\tilde{E}_{m_{i}}=\langle m_{i}|\tilde{H}_{0,n}|m_{i}\rangle$, and $\tilde{E}_{i}=\langle gn|\tilde{H}_{0,n}|gn\rangle=0$. For an odd $l$, $\Omega_{\rm eff}^{(l)}$ represents the effective coupling strength for connecting $|gn\rangle$ and $|hn\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|hn\rangle$. For an even $l$, it represents the effective coupling strength for connecting $|gn\rangle$ and $|e(n+1)\rangle$ through the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|e(n+1)\rangle\cdots\rightarrow|e(n+1)\rangle$. Similarly, the coefficient of the effective Hamiltonian (7) is the real part of the summation of all round trips about $|gn\rangle$
$\langle gn|H_{n}^{\rm eff}|gn\rangle=$ ${\rm Re}\left[-\frac{\Omega^{2}}{\tilde{\Delta}}\sum_{k=0}^{\infty}\left(\frac{g^{2}A_{n}^{2}}{\tilde{\Delta}\tilde{\delta}}\right)^{k}\right]$ (14)
$=$ ${\rm Re}\left[-\frac{\Omega}{\tilde{\Delta}}\sum_{k=0}^{\infty}\Omega_{\rm eff}^{(2k+1)}\right].$
In parallel to appendix A, Eqs. (11)-(14) constitute a straightforward and intuitive method to obtain the effective Lindblad operators (10a) and (10b) and the effective Hamiltonian (9) by taking the decay amplitude $\sqrt{\gamma_{hg}}$ and $\sqrt{\gamma_{eg}}$, the Rabi frequency $\Omega$, and the effective splittings $\tilde{\Delta}$ and $\tilde{\delta}$ of all the subspaces into account. Alternatively, they give rise to the same effective master equation (8).
@@
[…]
The justification of our charging scheme begins with a battery of a uniform energy ladder configuration, i.e., $A_{n}=1$ that is popular in literature Seah et al. (2021); Chen et al. (2022); Mitchison et al. (2021), resulting in an $n$-independent effective transition rate across all the subspaces. Substituting the optimized condition (16) to Eq. (15), it is found that the optimized rate reads